OpenStudy (anonymous):

Consider the line y=4x-1 Find the equations of the line that is perpendicular & parallel to this line and passes through the points (-2,3)

OpenStudy (ranga):

Do you know how to find the slope of the line y = 4x - 1 ?

OpenStudy (anonymous):

the slope would be 4/1 wouldnt it?

OpenStudy (ranga):

Yes, but just 4. Because the general equation of a line is: y = mx + b where m is the slope. So if you compare y = mx + b to y = 4x - 1 then you know m = 4. Now the slope of the perpendicular line will be just the negative reciprocal of the slope of this line. So slope of perpendicular line is: -1/4 So the equation of the perpendicular line is: y = -1/4x + b You will have to solve for b. They tell you the line passes through (-2, 3) Put x = -2 and y = 3 and solve for b.

OpenStudy (anonymous):